OpenStudy (rebecca1588):

Metal + Fan The table shows values for points on the graph of a linear function. x −2 1 2 4 y 13 1 −3 −11 What is the slope of the graph of this linear function? ____________

OpenStudy (4everaddicted2anime):

To find the slope you need the following formula \[\frac{ y_2-y_1 }{ x_2-x_1 }\] Then you need to pick two points and plug those into the formula. I'll use the points (-2,13) and (1,1)\[\frac{ 1-13 }{ 1--2 }\] After simplifying you should get\[\frac{ -12 }{ 3 }=-4\] The slope will be -4

OpenStudy (rebecca1588):

Thank-chuuuuuuu!!!!

OpenStudy (4everaddicted2anime):

yw